### v1

# Quantum amplitude estimation from classical signal processing

Abstract
We demonstrate that the problem of amplitude estimation, a core subroutine used in many quantum algorithms, can be mapped directly to a problem in signal processing called direction of arrival (DOA) estimation. The DOA task is to determine the direction of arrival of an incoming wave with the fewest possible measurements. The connection between amplitude estimation and DOA allows us to make use of the vast amount of signal processing algorithms to post-process the measurements of the Grover iterator at predefined depths. Using an off-the-shelf DOA algorithm called ESPRIT together with a compressed-sensing based sampling approach, we create a phase-estimation free, parallel quantum amplitude estimation (QAE) algorithm with a total query complexity of $\sim 4.9/\varepsilon$ and a parallel query complexity of $\sim 0.40/\varepsilon$ at 95% confidence. This performance is a factor of $\times1$ and 14×14$\times$14 × improvement over Rall and Fuller [Quantum 7, 937 (2023)], for worst-case complexity, which to our knowledge is the best published result for amplitude estimation. The approach presented here provides a simple, robust, parallel method to performing QAE, with many possible avenues for improvement borrowing ideas from the wealth of literature in classical signal processing.
I Introduction
Amplitude Estimation (AE) [1] is a fundamental quantum algorithm with many applications. For example, it provides a quadratic speedup in Monte Carlo methods [2], giving speedups to problems in the financial sector [3, 4, 5, 6]. AE is also a subroutine used to improve the complexity of algorithms that must estimate, for example, overlaps of states at the end of the quantum linear system algorithm [7, 8].
AE was first introduced as a combination of Grover search [9] and Quantum Phase Estimation (QPE) [10]. However, it was long conjectured that QPE was not necessary since it typically provides exponential speedup, while AE only provides a quadratic speedup over classical algorithms. This was proven true when Suzuki et al., showed that using Grover’s algorithm combined with classical maximum likelihood estimation based post-processing achieved the optimal scaling without requiring QPE [11]. This approach has been improved [12, 13, 14] with the result that achieves the best known query complexity, to our knowledge, being the one based on quantum signal processing [15].
Quantum algorithms for AE without QPE take measurements of the quantum state at different values of $n$, the number of applications of the Grover iterator, and use classical post-processing either at the end [13, 16, 15] or iteratively [12, 14] to determine at what $n$ to take samples next. The downside to the iterative approaches, as was pointed out in Refs. [11, 13] is that one has to switch between quantum and classical repetitions in a serial manner, which may be undesirable in practice.
Here, we demonstrate a non-iterative AE algorithm with the benefit that the number of iterations is known from the outset, so every sample can be done in parallel. In addition, the classical post-processing is a) robust to noise, allowing one to take very few samples to achieve low overall query complexity, and b) is classically efficient, not affecting the overall algorithm complexity up to log factors. Finally, in comparison to other similarly performant QAE algorithms, our approach does not require quantum signal processing (QSP) which simplifies the fault-tolerant circuits required to implemnt it. To achieve these results, we draw from ideas in (classical) signal processing, in particular from algorithms used in determining the Direction Of Arrival (DOA) of an incoming signal.
In DOA a set of sensors is placed at certain positions in space to detect an incoming signal from an unknown position [17, 18, 19, 20, 21]. The measurement data is then used to determine the angle of arrival of the incoming signal relative to the position of the sensors. This problem has been well-studied as it has numerous applications in fields such as radar, sonar, and wireless communication. There are various algorithms for doing this, for example the MUltiple SIgnal Classification (MUSIC) algorithm [22] or Estimation of Signal Parameters via Rotational Invariance Techniques (ESPRIT) [23] are from a class of subspace methods, but there are many more variants and sampling techniques [20].
The number of sensors and the spacing determines the accuracy or resolution with which one could estimate the DOA. Minimizing the number of sensors while simultaneously achieving high accuracy is desired for high-performance. It was discovered that using compressed-sensing methods allowed one to create what were called virtual arrays from sparse arrays, where the effective sensor spacing was greater than the number of actual sensors.
One example are coprime arrays, where one combines two uniform samplers with sample spacings $MT$ and $NT$ where $M$ and $N$ are coprime integers and $T$ has dimension of space or time. This allows one to generate $\mathcal{O}(MN)$ sample locations using only $\mathcal{O}(M+N)$ physical samples [24]. This can be further improved using multiple level nested arrays to achieve $\mathcal{O}(N^{2q})$ virtual sensors for some integer $q$ using just $\mathcal{O}(N)$ physical sensors [25].
It turns out that there is a nearly one-to-one correspondence between the DOA estimation problem and the amplitude estimation problem. This allows us to convert quantum measurements into a signal that can be post-processed using these DOA algorithms to estimate the amplitude. Our results demonstrate just one type of sensor spacing approach, the $2q$ array [25] and one post-processing algorithm ESPRIT [23]. We note that there are a wide number of variations of different sampling strategies and post-processing algorithms that could be further explored to optimize this approach to particular quantum algorithm and application needs [20].
II Classical signal processing based AE
The setting is as follows: suppose we have access to a unitary $U$ such that
$U\ket{0^{l}}$ $=\cos\theta\ket{x,0}+\sin\theta\ket{x^{\prime},1}$ (1)
$=\sqrt{1-a^{2}}\ket{x,0}+a\ket{x^{\prime},1}$
for some $x,x^{\prime}$ (states on $l-1$ qubits) and unknown $\theta\in[0,\pi/2]$. We wish to design an algorithm that finds $\theta$ (or the amplitude $a=\sin\theta$) up to an additive error $\varepsilon>0$. Classically, one would simply measure the state and conditioned on measuring a |$\ket{0}$ or |$\ket{1}$ in the final qubit and estimate $\theta$ using $\mathcal{O}(1/\varepsilon^{2})$ samples, or applications of the unitary $U$. Using AE, this can be improved to the optimal scaling that instead requires only $\mathcal{O}(1/\varepsilon)$ samples or applications of the unitary $U$, giving a quadratic speedup for quantum over classical.
To achieve the optimal scaling, let $R_{0}$ be the reflection in |$\ket{0^{l}}$, that is $R_{0}\ket{x}=-\ket{x}$ for $x\neq 0^{l}$ and $R_{0}\ket{0^{l}}=\ket{0^{l}}$. Let $S_{0}$ be the reflection in |$\ket{0}$ in the last qubit, that is $S_{0}\ket{x,0}=\ket{x,0}$ and $S_{0}\ket{x,1}=-\ket{x,1}$ for all $x$. We can define the so-called Grover operator $G:=UR_{0}U^{-1}S_{0}$ which has the following property: after $n$ applications the state becomes
|$\ket{\phi_{n}}$ $:=G^{n}U\ket{0^{l}}$ (2)
$=\cos((2n+1)\theta)\ket{x,0}+\sin((2n+1)\theta)\ket{x^{\prime},1}.$
When we measure the last qubit of the state |$\ket{\phi_{n}}$ in the computational basis we obtain |$\ket{0}$ or |$\ket{1}$ with probabilities
$p_{0}(n)$ $:=\cos^{2}((2n+1)\theta)$ (3a)
$p_{1}(n)$ $:=\sin^{2}((2n+1)\theta).$ (3b)
When measuring |$\ket{\phi_{n}}$ in the $X$-basis (apply the Hadamard gate and then measure), we obtain |$\ket{0}$ or |$\ket{1}$ with probabilities
$p_{0}^{X}(n)$ $:=\frac{1}{2}(1+\sin((2n+1)2\theta))$ (4a)
$p_{1}^{X}(n)$ $:=\frac{1}{2}(1-\sin((2n+1)2\theta)).$ (4b)
This allows us to construct the estimator
$\frac{p_{0}^{X}(n)-p_{1}^{X}(n)}{p_{0}(n)-p_{1}(n)}=\tan((2n+1)2\theta).$ (5)
Taking the arctan of the expression above we obtain an approximation of $(2n+1)2\theta$ from which we can form the exponential
$y_{n}=e^{i(2n+1)2\theta}=e^{in4\theta+2i\theta}.$ (6)
The need to take the arctan instead of just the arccos or arcsin of Eq. (4) is needed to span the entire range of $20\leq\theta\leq\pi/$ 2.
The estimation of $y_{n}$ is not exact. The probabilities that we estimate from Eqs. (3) and (4) have binomial sampling noise. The goal of AE is to estimate $\theta$ using as few measurements as possible. In the next section, we describe a classical signal processing approach that uses the estimator defined in Eq. (6), mimicking almost exactly the same type of estimator one obtains using physical sensors that are attempting to estimate the direction of arrival of an unknown signal using noisy measurements.
II.1 Subspace Based Direction Of Arrival Estimation
Suppose a linear array of physical sensors placed at positions $x_{1},\dots,x_{M}$ and suppose that there is one source providing an incoming signal with arrival angle $\bar{\omega}\in(-\pi/2,\pi/2)$ shown schematically in Fig. 1. A sensor at position $x_{n}$ obtains a measurement of
$y_{n}^{\prime}=e^{i2\pi(n-1)d\sin\bar{\omega}/\lambda}+\epsilon_{n}=e^{i\omega x_{n}}+\epsilon_{n}=y_{n}+\epsilon_{n},$ (7)
where we have identified a unit distance $x_{n}=(n-1)$ and angle $\omega=2\pi d\sin{\bar{\omega}}/\lambda$ for an incoming signal with wavelength $\lambda$. The parameter $y_{n}$ here is distinct from that defined in Eq. (6), but we purposely use the same notation as it turns out they are equivalent from the perspective of DOA estimation. The parameter $\epsilon_{n}$ denotes an error term due to imperfect measurements. The goal of DOA estimation, is to determine the incident angle of the incoming signal $\bar{\omega}$ using the fewest possible sensors. Looking at Eq. (6), we can view it as an incoming signal with angle of incidence $\omega=4\theta$ at position $n$. This is the main connection between AE and DOA estimation that we will use. The main difference is how we obtain noisy measurements of $y_{n}$.
Figure 1: Figure depicting the direction of arrival estimation geometry for a uniform linear array. Here a linear array of $M$ sensors is arrayed and the phase of the incident wavefront depends upon its angle of arrival.
Consider now the covariance matrix, $\boldsymbol{R}$, of the estimated signal vector $\boldsymbol{y}^{\prime}=(y_{1}^{\prime},\ldots,y_{M}^{\prime})^{T}$
$\boldsymbol{R}=E[\boldsymbol{y}^{\prime}\boldsymbol{y}^{\prime\dagger}]=\boldsymbol{y}\boldsymbol{y}^{\dagger}+\sigma^{2}\boldsymbol{I},$ (8)
where $\sigma^{2}=E[\epsilon_{j},\epsilon_{j}]$ is the variance of the noise which is assumed to be Gaussian and uncorrelated, and $\boldsymbol{I}$ is the identity matrix. One of the key insight of subspace based DOA estimation algorithms, is that the covariance matrix can be decomposed into signal and noise subspaces, which can lead to accurate estimation of the DOA in the presence of noisy estimates. Consider the eigendecomposition of $\boldsymbol{R}$
$\boldsymbol{R}\boldsymbol{u}_{n}=\left(\boldsymbol{y}\boldsymbol{y}^{\dagger}+\sigma^{2}\boldsymbol{I}\right)\boldsymbol{u}_{n}=\lambda_{n}\boldsymbol{u}_{n}.$ (9)
The eigenspace of the covariance matrix $\boldsymbol{R}$ can be partitioned into a signal subspace and a noise subspace. For AE there is only one signal, so the signal subspace has dimension 1 with eigenvalue $\lambda_{s}=\sigma_{s}^{2}+\sigma^{2}$ and eigenvector $\boldsymbol{u}_{1}$, while the noise subspace is size $M-1$ with eigenvalues $\lambda_{n}=\sigma^{2}$ and eigenvectors $\boldsymbol{u}_{n\neq 1}$.
It turns out that the eigenvectors associated with the noise subspace are orthogonal to the signal vector $\boldsymbol{y}$ [26]. That is, one can show that
$\boldsymbol{y}^{\dagger}\cdot\boldsymbol{u}_{j}=0\quad j\neq 1.$ (10)
This approach was one of the first subspace based approaches to DOA estimation, called the MUSIC algorithm. It estimates the angle of arrival of a signal, $\omega^{\prime}$, by finding angles where the signal subspace is orthogonal to the noise subspace [26]. The MUSIC algorithm performs well in practice, but it is overkill for what is required for AE. It requires one to perform a full eigendecomposition of the $M\times M$ covariance matrix $\boldsymbol{R}$ which has complexity $\mathcal{O}(M^{3})$ and also requires one to compute Eq. (10) for many different possible angles $\omega^{\prime}$ to achieve good performance.
Standard DOA can handle multiple sources, but this is not necessary for AE. However, we want to note that having the ability to handle multiple sources might be useful to optimize a core part in the quantum mean estimation algorithm of [27]. In their approach, the Grover operator that is used is not just 2-dimensional (in AE the Grover operator is essentially a rotation on a plane), but can be high-dimensional. In signal processing terms, this corresponds to having multiple sources of direction of arrival.
II.2 ESPRIT Approach to DOA Estimation
Just like MUSIC, ESPRIT [23] is an algorithm that uses the measurements $\boldsymbol{y}^{\prime}$ to estimate the unknown angle $\omega$. The difference is that instead of estimating the incoming spectrum using the noise eigenspace as shown in Eq. (10), it just estimates the incoming DOA directly from the signal subspace. This allows a much more computationally efficient approach to DOA estimation. We outline this in Algorithm 1 and refer the reader to Refs. [23, 28, 21] for further details.
Algorithm 1 ESPRIT
0:  Measurements of a signal $y(x)=e^{i\omega x}$ on a uniformly spaced array of sensors.
0:  An approximation of $\omega$ up to error $O(1/M)$ where $M$ is the size of the array of sensors.
1:  Form the Toeplitz matrix [28] $\boldsymbol{R}$ with first row the vector of measurements $\boldsymbol{y}^{T}\in\mathbb{C}^{M}$ and first column the vector of measurements $\boldsymbol{y}^{*}\in\mathbb{C}^{M}$ and compute it’s singular value decomposition $\boldsymbol{R}=\boldsymbol{U}\boldsymbol{\Sigma}\boldsymbol{V}$.
2:  Form the matrix $\boldsymbol{S}$ from the first 2 columns of $\boldsymbol{U}$, the matrix $\boldsymbol{S}_{1}$ from the first $M-1$ rows of $\boldsymbol{S}$, and the matrix $\boldsymbol{S}_{2}$ from the last $M-1$ rows of $\boldsymbol{S}$.
3:  Compute the $2\times 2$ matrix $\boldsymbol{P}=\boldsymbol{S}_{1}^{-1}\boldsymbol{S}_{2}$ where $\boldsymbol{S}_{1}^{-1}$ is the pseudo-inverse.
4:  Output the phase of the first eigenvalue of $\boldsymbol{P}$.
For AE, we don’t need to do the full singular value decomposition of the matrix $\boldsymbol{R}$ because we know from the outset that there is just a single incoming signal. Instead, we just need the top two eigenvectors to form the matrices $\boldsymbol{S}_{1}$ and $\boldsymbol{S}_{2}$. Because the matrix is Toeplitz, we can use the Lanczos algorithm [29] together with Fourier-transform based algorithms for computing matrix-vector products of Toeplitz matrices to obtain these two eigenvectors in time $O(M\log(M))$ [30]. Computating the pseudo-inverse of $\boldsymbol{S}_{1}$ is likewise fast, taking time $\mathcal{O}(M)$, since the matrix is of size $(M-1)\times 2$. Therefore the classical complexity of using ESPRIT for AE is $\mathcal{O}(M\log(M))$, logarithmically worse than the lower-bound query complexity of AE, but insignificant in practice. We find that the post-processing for the largest matrix we numerically simulate in this manuscript with $M=215177$ and targets a 95%-percentile error rate of $9.85\times 10^{-6}$, takes 8 seconds on a laptop with the current implementation and probably can be optimized much further.
II.3 Physical and virtual arrays
The previous descriptions, assumed a uniform array of sensors, denoted as a uniform linear array (ULA). However, this can be prohibitively costly as the required number of sensors can be large for good performance. A major advancement in this field was the development of the theory of sparse array processing [20] where one places sensors at only a few sub-sampled locations of the ULA which is called the physical array. The virtual array concept is the idea that one can use the data from the physical array to get signal measurements at positions where there is no physical sensor, by combining the measurements of the signal from the physical array in a certain way.
The ESPRIT algorithm, described previously, uses measurements of the incoming signal at physical positions which are evenly spaced. The virtual array concept allows us to obtain such measurements by using fewer physical sensors. As an example, suppose we place the sensors at physical locations $x=(x_{1},x_{2},\dots,x_{n})$, which is a subset of sensor positions needed for the uniform linear array. Consider the vector of signals, $\boldsymbol{y}$, at those positions $y_{j}=e^{i\omega x_{j}}$. Compute the outer product of $y$ with itself, to obtain the following matrix
$(\boldsymbol{y}\boldsymbol{y}^{\dagger})_{ij}=e^{i\omega(x_{i}-x_{j})}.$
This is the value of the signal value at the physical location $x_{i}-x_{j}$. Suppose we have noisy measurements $y_{j}^{\prime}$ of $y_{j}$ at locations $x_{j}$. We can use the value $(\boldsymbol{y}^{\prime}\boldsymbol{y}^{\prime T})_{ij}$ to estimate the signal $\boldsymbol{y}$ at location $x_{i}-x_{j}$ even though we never physically measured the signal at that location. This technique allows us to obtain many more (virtual) measurements than the number of physical sensors. This approach is what allows us to achieve the optimal $\varepsilon\sim 1/N$ scaling for AE.
Given a vector $x\in\mathbb{R}^{n}$ that represents a linear array of physical sensors where sensor $i$ is placed at position $x_{i}$, we can define for integer $q\geq 1$ its virtual array by taking repeated outer products $q$ times.
Definition II.1 (Virtual array [31, 32]).
Let $x=(x_{1},\dots,x_{n})\in\mathbb{R}^{n}$ and $q\geq 1$ an integer. The $2q$-th order virtual array corresponding to $x$ is given by
$S_{q}:=\{\sum_{i=1}^{q}x_{k_{i}}-\sum_{i=q+1}^{2q}x_{k_{i}}\colon k_{i}\in[n]\}.$ (11)
It is important to minimize the size of the physical array while maximizing the size of the virtual array. As we will see in the next section, the elements of the physical array $x_{j}$ corresponds to the number of Grover operators we apply.
We can create long virtual arrays using the following theorem on choosing the physical sensor locations [25].
Theorem II.2.
Let $q,N_{1},N_{2},\dots,N_{2q}$ be positive integers and consider the following sets for $11\leq i\leq 2q-$ 1
$\{n\prod_{k=0}^{i-1}N_{k}\colon n=1,2,\dots,N_{i}-1\},$ (12)
and the $2q$-th set is given by $\{n\prod_{k=0}^{2q-1}N_{k}\colon n=1,2,\dots,N_{2q}\}$ where $N_{0}:=1$. Let the physical array be the union of all these sets. Then the $2q$-th order virtual array corresponding to this physical array contains a ULA of size
$12\prod_{k=0}^{2q}N_{k}-$ 1 . (13)
As an example, we choose $N_{1}=N_{2}=\dots=N_{2q}=2$ so that the union of all the sets in equation (12) is equal to $(2^{j})_{j\in[2q]}$ and the corresponding virtual array has size $2^{2q+1}$. Fig. 2 shows an example where $q=2$. The physical array is given by $\{1,2,4,8\}$ which is shown by the green stars (also including the location 0). The second-order virtual array is the next level given by the dark blue dots, this is obtained by taking all the possible differences of pairs in the physical array. The 4-th order virtual array is given by the cyan triangles and it contains a ULA of size 31. Note that we only ever take measurements at the green dots, whose size (the number of locations) is significantly smaller than the virtual array (cyan dots) which is quantified by the above Theorem.
Once the virtual array is created using this method, we create what is known as a spatially smoothed version of the covariance matrix $\boldsymbol{R}$ from the virtual array using the method of Ref. [28]. This results in a Toeplitz covariance matrix that allows for efficient classical post-processing as remarked in the previous section.
Taking samples at depths of powers of two is reminiscent of the Maximum Likelihood Amplitude Estimation (MLAE) approach to AE [11]. The main difference here is the classical post-processing, which provides far better results than MLAE. This was well-known for some time in the signal processing literature [21]. We used powers of two to determine scaling factors of our approach in Sec. IV.1, but note that there is a great deal of flexibility here depending on the use-case. We highlight some alternative schedules when discussing the impact of circuit noise in Sec. IV.3.
Figure 2: Physical and virtual locations for $q=2$ and $N_{1}=N_{2}=N_{3}=N_{4}=2$. Note that the $2q$-th order (in this case 4th order) virtual array contains a large ULA.
III The AE algorithm
We now show how these DOA algorithms outlined in the previous section can be used for AE. Upon examination of Eq. (6) and Eq. (7), we see that the measurement model in the two approaches are identical and the DOA estimation methods are blind to the extra phase factor $e^{2i\theta}$ in $y_{n}$, meaning it will extract just $4\theta$ from measurements of $y_{n}$.
In classical signal processing terms, we can obtain an estimate of the signal $y_{n}=e^{in4\theta+2i\theta}$ by placing a physical sensor at location $n$. Quantumly, we can obtain a estimate of the signal $y_{n}$ by taking repeated measurements of |$\ket{\phi_{n}}$ in the $Z$ and $X$ basis. As far as the DOA algorithms are concerned it does not matter how the measurements are obtained.
The next step is to determine at what depths we take measurements. The DOA algorithms expects measurements on a uniform linear array of length $M$, that is measurements of $y_{n}$ for $n\in[M]$. One may obtain this by measuring |$\ket{\phi_{n}}$ at depths all $n\in[M]$, with a query complexity of $\sum_{n=1}^{M}n=O(M^{2})$. This does not provide us with the expected quantum scaling $O(1/M)$ of the error, but rather the scaling is $O(1/\sqrt{M})$ in this case. We avoid this problem by taking samples from a far smaller set of depths and use the virtual array concept to obtain (virtual) measurements at all depths $n\in[M]$.
More precisely, we apply Theorem (II.2) and use the physical array given by the sequence $D=(2^{j})_{j\in[2q]}$. The query complexity of taking measurements of |$\ket{\phi_{n}}$ for all $n\in D$ is $O(2^{2q+1})$, while the size of the ULA inside the virtual array is $2^{2q+1}$ by Theorem (II.2). Since the DOA algorithms can estimate the incident angle with accuracy $O(1/M)$ if $M$ is the length of the ULA, this implies immediately the quantum scaling that we were looking for. A summary of the full AE algorithm is given in Algorithm 2.
Algorithm 2 Amplitude Estimation from Classical Signal Processing (csAE)
0:  Oracle $U$ such that $U\ket{0^{l}}=\cos\theta\ket{x,0}+\sin\theta\ket{x^{\prime},1}$ and error rate $\varepsilon>0$.
0:  Approximation of $a=\cos(\theta)$ to additive error $\varepsilon$.
1:  Let $D$ be a set of depths such that the $2q$-th order virtual array has size $M\approx 1/\varepsilon$.
2:  Measure the state |$\ket{\phi_{n}}$ for $n\in D$ in the $Z$ and $X$ basis and form the “signal” vector $y_{n}=e^{i\omega_{n}}$ where $\omega_{n}$ is the arctan of the expression in (5) and where the $p_{i}(n)$ and $p_{i}^{X}(n)$ are replaced by their empirical values (from the measurements).
3:  Form the signal vector $\boldsymbol{y}\in\mathbb{C}^{M}$ by computing the signal values at the virtual locations corresponding to positive virtual positions.
4:  Use the ESPRIT (or any classical DOA) algorithm with $\boldsymbol{y}$ as input to extract the angle $\hat{\omega}$.
IV Numerics
Figure 3: Estimation error as a function of query complexity (top) and parallel complexity (bottom) for amplitude $a=0.5$. For csAE we simulated 500 runs giving a range of error rates (given by the vertical dots). The green triangles denote the 95-th percentile of the errors and the green fit line is used to compute the scaling factor. For chebAE we also simulated 500 runs, for a target $\varepsilon$ given by the results from csAE. The cloud of points from chebAE arise because of the iterative nature. The red and cyan stars and fit lines denote the average and worst case complexity respectively for chebAE for the 95-th percentile.
Refer to captionRefer to caption
Figure 4: Scaling factor $C_{\textrm{opt}}$ as a function of amplitude for total (top) and parallel complexity (bottom) of csAE (green triangles) and chebAE (red stars) at $\delta=0.95$ confidence. The horizontal lines correspond to the worst scaling factor for the amplitudes tested and is the number we report as the overall scaling factor.
In this section we show the performance of our algorithm csAE and compare it against the state of the art Amplitude Estimation algorithm which we refer to as chebAE [15]. The code for reproducing the numerics and plots in this paper can be found on https://github.com/unitaryfund/csAE.
We compare the number of oracles queries required by the algorithm to estimate the amplitude $a$ defined in Eq. (1). The metric we compare is the estimation error
$\varepsilon=|a-\hat{a}|$ (14)
where $\hat{a}$ is the estimated amplitude error rates of the algorithm from different runs. More formally, we also define a confidence level of our estimate. That is we estimate
$\textnormal{Pr}[|a-\hat{a}|<\varepsilon]\geq\delta$ (15)
where $\delta$ is the confidence level.
The query complexity is the number of times we use the operator $U$ (see  (1)) divided by two. Since the Grover operator $G$ consists of two calls to $U$, the query complexity will be the number of times we use $G$ with an additional number of oracle calls to $U$ at depth 0 divided by two. In chebAE [15] the number of oracle queries is defined as the number of times we apply the Grover operator $G$ ignoring the calls to $U$ at depth 0. This won’t affect constant factors in any meaningful way (we are penalizing our algorithm slightly), but the way we account for the queries seems to be fairer. We compare the two approaches ability to achieve a target $\varepsilon$ and $\delta$ on two different metrics: 1) query complexity and 2) parallel query complexity. By parallel query complexity we mean the maximum depth oracle query.
To get a better intuition of the performance of the algorithm, in Fig. 3, we provide results when estimating the amplitude $a=0.5$ for 500 random trials where the only source of randomization is the binomial sampling error when estimating the probability. We use a physical array with parameters $N_{1}=N_{2}=\cdots N_{2q}=2$ as defined in Eq. (12) for $q\in\{3,4,\ldots,8\}$. For csAE, the input is the physical sampling array, denoting the fixed query depth, as well as the number of samples to take at each depth. This results in a deterministic schedule and yields the vertical colored dots representing the estimation errors of the algorithm from different runs. The green triangles represent the $\delta=0.95$ confidence level of those runs for different depth circuits. The green line is a fit line to these 95% confidence levels.
The number of samples to take at each depth is user-defined. We found that using a fixed number of shots was sub-optimal. The algorithm yielded better constant factor complexity when taking more shots at shorter depths and fewer shots at deeper circuit depths. As a heuristic, we used the following schedule
$N_{\mathrm{shots}}(n)=\lceil K(\log_{2}(n_{\mathrm{max}}/n)+1)\rceil,$ (16)
where $n_{\mathrm{max}}$ is the maximum query depth, $n$ is the current query depth, and $K$ is a constant that we chose to be $K=1.3$ for $\delta\leq 0.95$ and $K=1.8$ for $\delta=0.99$. This was determined heuristically, so it is possible that there exist better schedules $N_{\mathrm{shots}}(n)$. We note that it is also possible to have a non-deterministic schedule that utilizes information from the the top-two eigenvalues of the matrix $\boldsymbol{P}$ of Algorithm 1. When these eigenvalues are close to the remaining eigenvalues the signal and noise subspaces are not well separated. We believe that using this information could further improve the constant factor scaling, but at the cost of reducing parallelizability. We leave this to future work.
We directly compare our results to the performance of chebAE in the same figure. In the case of chebAE, the estimation error and confidence-level are inputs to the algorithm while the query complexity is a random variable determined at runtime. This results in the cloud of points observed on the plot, which is the result of each independent run of chebAE. Because the query complexity is a random variable, we report both the $\delta=0.95$ confidence level for the average query complexity and worst case query complexity as the red and cyan points respectively. The red and cyan lines are fit lines to these 95% confidence levels. In addition to the determinstic schedule benefit of csAE, we also note that the worst-case error rate is typically more constrained compared to chebAE, which can fail with much larger error rates for the parameters chosen.
IV.1 Fits
Table 1: Fitting constants for various confidence levels, along with the associated standard error obtained from the weighted-least-squares fit. Overall, csAE is comparable to the average case complexity of chebAE in terms of total query complexity, but tends to outperform in terms of worst-case complexity. In addition, csAE is highly parallelizable resulting in vast out-performance for parallel query complexity.
$99\%$ total $99\%$ parallel $95\%$ total $95\%$ parallel $68\%$ total $68\%$ parallel
csAE $8.5\pm 1.6$ $0.6\pm 0.1$ $4.9\pm 0.6$ $0.40\pm 0.05$ $2.0\pm 0.1$ $0.162\pm 0.008$
chebAE (worst) $8.2\pm 0.5$ $8.0\pm 0.3$ $5.7\pm 0.2$ $5.4\pm 0.1$ $2.52\pm 0.08$ $2.23\pm 0.09$
chebAE (ave) $5.8\pm 0.4$ $5.6\pm 0.3$ $4.0\pm 0.1$ $3.77\pm 0.09$ $1.55\pm 0.08$ $1.44\pm 0.05$
To estimate the scaling factor, we perform a weighted least-squares fit to the function $N=C/\varepsilon_{\delta}+b$ ($N$ number of queries and $\varepsilon_{\delta}$ error rate at a given confidence $\delta$), where the weighting factor is given by $\varepsilon_{\delta}$. We get an estimate of the optimal parameter $C_{\mathrm{opt}}$ by minimizing the weighted residuals in the 2-norm: write $r_{i}:=N_{i}-C/\varepsilon_{\delta,i}-b$ for the residual of observation $i$ and $w_{i}=\varepsilon_{\delta,i}$ the corresponding weight, then
$C_{\mathrm{opt}}:=\min_{C}\sum_{i}w_{i}r_{i}^{2}.$ (17)
The performance of the algorithm depends on the value of the amplitude $a$ to be estimated. Since $a$ is an unknown parameter, we compare the performance of csAE for amplitudes in the range $[0.1,0.2,\ldots,0.9]$. For each amplitude, we perform the same analysis shown in Fig. 3 and compute the estimation error $\varepsilon_{\delta}$ at a given confidence-level $\delta$. We compute the weighted-least squares fit using both average and worse-case query complexities for each amplitude to compute the scaling factor $C_{\textrm{opt}}$.
We plot the results of these fits for $\delta=0.95$ confidence in Fig. 4 for both total query complexity and parallel query complexity using the worst-case query complexity. We report the constant factor as the maximum constant factor as a function of amplitude. We find that csAE generally performs better for larger amplitudes, while chebAE has worst-case complexity at $a=0.5$ as previously reported [15].
In Tab. 1 we provide a summary of the constant factor scaling comparing csAE to chebAE for different confidence levels and for both average and worst-case query complexity for chebAE. Overall, we find that our algorithm slightly under-performs chebAE for the average case, but over-performs for the worst case, however many of these are within the error bars.
While csAE does not significantly lower the constant factor in terms of query complexity compared to chebAE, the implementation is much simpler. As with the IQAE implementation [14], the csAE approach makes use of the standard Grover iterator and do not require QSP. This implies that the circuit-level implementation is substantially simpler than for chebAE, not requiring controlled phase rotations and circuit synthesis, which adds a logarithmic factor to the required fault-tolerant circuit depth [33]. Therefore, our approach yields the best of both worlds: simple circuit construction, together with state-of-the-art query complexity. As we discuss next, our approach also offers a highly parallel implementation, something not available to either chebAE or IQAE.
IV.2 Parallelizability
Figure 5: Scaling of the constant factor for the parallel query complexity as we increase the number of quantum computers. This is for the amplitude $0.5$ and the depth schedule is $(2^{j})_{j\in[2q]}$ for $q=5$ and shots schedule $[15,13,12,11,10,8,7,6,4,3,2]$. This gives a total query complexity of $6,417$ and results in an error rate of $5.6\times 10^{-4}$ (with $95\%$ confidence). Because shots at lower $q$ values require fewer oracles calls, they can be parallelized across fewer computers resulting in the saturation near 12 QPUs.
Our algorithm is fully parallelizable, because the quantum circuits at which we take samples from is known beforehand. Here we look at the parallel query complexity for a given number of quantum computers. The idea is to divide the total number of queries as evenly as possible among the QPUs. The parallel query complexity is then simply the largest query complexity among the QPUs. We can find this even division using a greedy approach: sort the circuits from which we take samples in decreasing Grover power and distribute the circuits evenly among the QPUs starting with the highest Grover power. In Fig. 5 we plot the constant factor $C=N\cdot\varepsilon$ for the parallel query complexity given a number of QPUs. One can see that the constant factor saturates at some point (12 QPUs in this case), because the largest Grover power has to be run on some QPU, hence providing a lower bound on the parallel query complexity.
In quantum computer architectures where different Grover powers can be executed on the same QPU, the parallelizability of AE also allows for the application of the technique known as Hamming weight phasing (HWP) which efficiently applies repeated arbitrary single-qubit rotations by the same angle [34, 35]. When $m$ repeated rotations by the same angle $R_{z}(\theta)$ are parallelizable, this technique reduces the number of rotations which need to be synthesized to $\lfloor\log_{2}m+1\rfloor$, at the cost of $m-1$ Toffoli gates and $m-1$ ancilla qubits. Because the Grover operator consists of repeated applications of a unitary $U$ (and $U^{\dagger}$), any single-qubit rotations in $U$ can therefore be applied using HWP across different powers of the Grover operator which have been parallelized on the QPU. The practical benefit from this approach depends on the specific form of $U$ and the total number of Grover powers that must be applied, but we highlight this ability to reduce the total number of single-qubit rotations across different Grover powers as this is not possible in iterative variants of AE.
IV.3 Noise Resilience
Here we study the ability of our approach to handle noise in the quantum circuit. This could arise from current quantum devices that do not have error correction capabilities or future fault-tolerant devices where one might be interested in understanding what size quantum error correcting code is required for a given application.
We study this by implementing a simple noise model. When estimating the probability amplitudes at a given query depth $n$ as defined in Eqs. (3) and (4), we shift the probability by a per-oracle noise parameter $\eta$. That is, for the $Z$ basis measurements, we shift the probability to
$p_{0}(n)$ $\to\eta_{n}p_{0}(n)+(1-\eta_{n})/2$ (18a)
$p_{1}(n)$ $\to\eta_{n}p_{1}(n)+(1-\eta_{n})/2,$ (18b)
where $\eta_{n}=(1-\eta)^{n}$ is the noise parameter at a given depth $n$. As the circuit depth gets deeper, the noise parameter $\eta_{n}$ gets smaller resulting in the probabilities tending towards the completely mixed state. We implement a similar model for the $X$ basis measurements.
Figure 6: Estimation error as a function of total query complexity for amplitude $a=0.5$ and confidence $\delta=0.95$ for given values of the per-oracle noise parameter $\eta$. As $\eta$ increases, the ability of the algorithm to estimate the amplitude saturates as one would expect.
Table 2: Total number of queries, maximum coherent query depth, and array parameters required to achieve a target estimation error of $\varepsilon_{\delta}=1\times 10^{-3}$ for a given noise parameter $\eta$ and confidence level $\delta$. A dashed line indicates that the noise was too high and the estimation error could not be brought below the target threshold. The results suggest that for optimal performance at this estimation error, one should target a per-oracle error rate of $\eta\leq 10^{-5}$, where one essentially achieves noise-free performance.
$99\%$ total $99\%$ parallel $95\%$ total $95\%$ parallel $68\%$ total $68\%$ parallel
$\eta=10^{-3}$- -- $89,453$ $4,374$ $6,807$ $512$
$\eta=10^{-4}$ $18,262$ $1,152$ $8,399$ $648$ $2,311$ $192$
$\eta\leq 10^{-5}$ $10,214$ $512$ $6,004$ $360$ $2,311$ $192$
We plot the estimation error at $\delta=0.95$ confidence level for $a=0.5$ in Fig. 6 for a range of $\eta$ values. One can see that as the noise level increases, the ability of the algorithm to estimate the amplitude decreases exactly as one would expect. To quantify the performance of the algorithm, we report the number of queries required to achieve a given confidence level with estimation error $\varepsilon_{\delta}=1\times 10^{-3}$ which is generally consistent with many algorithmic performance requirements. These values are reported in Tab. 2.
In addition to the query counts, we also found that varying the array yielded better results. The reason, can be seen by examining Fig. 6. The target estimation error of $\varepsilon_{\delta}=1\times 10^{-3}$ lies between two arrays when using the schedule that increases the array size by powers of 2. However, as shown in Theorem II.2 the $N_{k}$ values do not need to be equal to the value 2, which we chose when reporting the scaling parameter. By choosing other integers, one can obtain an array that lies in the midpoint near where the estimation error crosses the target amount.
We tuned the array parameter values and the number of shots at each depth until we found an array that achieved the target estimation error and confidence level with the fewest total queries. This tuning was manual, so it is likely one could find better parameter values with more work. We report the total and parallel query complexities Tab. 2. We report a detailed summary of the array parameters, query depths, and shot scheduled used for those arrays in App. A.
V Discussion and Future Work
We have shown that there is a direct correspondence between quantum amplitude estimation and the estimation of the direction of arrival of an incoming signal, a well-studied problem in classical signal-processing. This correspondence allowed us to provide a new AE algorithm that is fully parallel and opens up the possibility of using the vast literature on DOA estimation to further optimize the performance of this algorithm.
We provide numerical results demonstrating one particular DOA algorithm ESPRIT that is computationally efficient, and achieves what is to our knowledge the best known query complexity by a small factor and the best known parallel query complexity by over an order of magnitude over previous approaches.
We can quantify the practical impact of the AE algorithm introduced in this manuscript by considering the problem of financial derivative pricing [3, 4, 5, 6]. In Ref. [36] the authors calculate that quantum advantage in pricing an autocallable derivative contract to accuracy $2\times 10^{-3}$ with confidence $68\%$ requires a T-depth of $4.5\times 10^{7}$, T-count of $2.4\times 10^{9}$ and in order for the calculation to meet the classical target of $1$ second, a $45$ MHz logical clock rate. This calculation assumes the IQAE variant [14] is employed at $\varepsilon=10^{-3}$ and confidence $\delta=68\%$ such that the total number of oracle calls is $C/\varepsilon=5735$ [5]. From the third column of Table 1, we observe that the total number of oracle calls to achieve the same accuracy and confidence using csAE is $2000$, a 2.9$\sim$ 2.9$\times$ 2.9 × reduction in both T-depth and T-count. Moreover, because csAE allows for parallelization, assuming it can be fully harnessed across multiple QPUs, the deepest circuit requires $162$ oracle calls (sixth column of Table 1), such that the logical clock rate required for quantum advantage is in theory reduced from $45$ MHz to $1.3$ MHz, a factor of $35$x reduction. However, while this figure might be of consequence in a discussion of whether a QPU with the calculated specifications for advantage can be constructed, it assumes that we parallelize the quantum algorithm, but not the classical counterpart used for comparison, Monte Carlo. Because Monte Carlo can be parallelized in a straightforward manner, a fair comparison would allow for both quantum and classical methods to execute in a parallel fashion. From Fig. 5, we observe that the csAE runtime scales as $\sim 1/N$ for $N$ processors, precisely the same way a classical Monte Carlo simulation scales with the number of processors. This means that no further logical clock rate benefit arises from the parallelization of csAE if we allow the same parallelization to Monte Carlo. Nevertheless, csAE does provide the possibility of parallelization, enabling trade-offs between circuit width and depth when desirable, unlike other high-performing AE variants like IQAE and chebAE, overcoming an important limitation of AE methods when compared to classical Monte Carlo.
There are a number of avenues to further improve or generalize our results depending on specific use cases. DOA estimation algorithms typically assume that the noise contained in the estimated signal is Gaussian and uncorrelated. However, the estimator we use to convert the quantum measurements to the form needed for DOA estimated, given in Eq. (5), has noise arising from the ratio of two binomially-distributed noise terms, which is manifestly non-Gaussian. In addition the estimator is biased. Despite this, our numerical results using an “off the shelf” implementation of ESPRIT already beat the previous best published results for worst-case complexity. We leave for future work further optimizations that may lead to even better constant factors by taking into account the actual noise distribution. Results in the DOA literature suggest that this can lead to substantial improvements (see e.g. Ref. [37]). We further remark that the noise resilience of this approach may make this method well-suited to NISQ-based algorithms that require amplitude-estimation as well [38, 39].
We chose a particular sampling strategy that we refer to as the $2q$-array. This choice, however, is not limiting and in fact there is a great degree of flexibility. We demonstrated the one degree of freedom that $2q$ arrays have in that the schedule can be fine-tuned to meet a required error tolerance, as shown in Sec. IV.3. One can also choose different array geometries such as co-prime arrays [24], ruler arrays [40], fractal arrays [41], along with many other approaches [42, 43]. Different array choices can lead to different asymptotic complexity $\mathcal{O}(1/\varepsilon^{\alpha})$ for various values of $\alpha$ allowing one to achieve lower circuit depth at the cost of needing more samples. This can be useful for noisy processors where circuit depth is limited [38, 39, 16].
We hope that this identification of quantum amplitude estimation as a special case of the well-studied direction of arrival estimation problem opens the door to many further improvements and generalizations.
Acknowledgements.
We thank Peter Johnson for constructive feedback on this manuscript and Patrick Rall, Bryce Fuller and Stefan Woerner for discussions on amplitude estimation.
Appendix A Array Parameters
Here we report detailed values for the array parameters used to generate the results presented in Tab. 2.
Table 3: Parameter values for $\varepsilon_{0.99}\leq 10^{-3}$.
Simulation parameters $\eta=10^{-3}$; $K=1.5$
Array Parameters --
Query Depths --
Shots per Depth --
Total Queries --
Simulation parameters $\eta=10^{-4}$; $K=1.8$
Array Parameters $[3,3,2,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,3,6,9,18,36,72,144,288,576,1152]$
Shots per Depth $[24,22,20,18,17,15,13,11,9,8,6,4,2]$
Total queries $18,262$
Simulation parameters $\eta=10^{-5}$; $K=2.1$
Array Parameters $[2,2,2,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,4,8,16,32,64,128,256,512]$
Shots per Depth $[24,21,19,17,15,13,11,9,7,5,3]$
Total Queries $10,214$
Table 4: Parameter values for $\varepsilon_{0.95}\leq 10^{-3}$.
Simulation parameters $\eta=10^{-3}$; $K=1.1$
Array Parameters $[3,3,3,3,3,3,3,3]$
Query Depths $[0,1,2,3,6,9,18,27,54,81,162,243,486,729,1458,2187,4374]$
Shots per Depth $[31,29,27,26,24,22,20,18,17,15,13,11,9,8,6,4,2]$
Total Queries $89,453$
Simulation parameters $\eta=10^{-4}$; $K=1.1$
Array Parameters $[3,3,3,3,2,2,2,2]$
Query Depths $[0,1,2,3,6,9,18,27,54,81,162,324,648]$
Shots per Depth $[15,14,13,11,10,9,8,7,6,5,4,3,2]$
Total Queries $8,399$
Simulation parameters $\eta=10^{-5}$; $K=1.3$
Array Parameters $[6,5,3,2,2,2]$
Query Depths $[0,1,2,3,4,5,6,12,18,24,30,60,90,180,360]$
Shots per Depth $[20,19,17,16,15,13,12,11,10,8,7,6,4,3,2]$
Total Queries $6,004$
Table 5: Parameter values for $\varepsilon_{0.68}\leq 10^{-3}$.
Simulation parameters $\eta=10^{-3}$; $K=1.5$
Array Parameters $[2,2,2,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,4,8,16,32,64,128,256,512]$
Shots per Depth $[17,15,14,12,11,9,8,6,5,3,2]$
Total Queries $6,807$
Simulation parameters $\eta=10^{-4}$; $K=1.1$
Array Parameters $[3,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,3,6,12,24,48,96,192]$
Shots per Depth $[11,10,9,8,7,6,5,4,3,2]$
Total Queries $2,311$
Simulation parameters $\eta=10^{-5}$; $K=1.1$
Array Parameters $[3,2,2,2,2,2,2,2]$
Query Depths $[0,1,2,3,6,12,24,48,96,192]$
Shots per Depth $[11,10,9,8,7,6,5,4,3,2]$
Total Queries $2,311$